The price P for a watermelon varies directly with its
weight w.
a) Write a formula for this variation.
b) If the price of a 30-pound watermelon is $4.20, then
what is the price of a 20-pound watermelon?

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

4.2/30=x/20
30x=84
x=$2.80

Answer by stanbon(75887)

The price P for a watermelon varies directly with its
weight w.
a) Write a formula for this variation.
P = kw
----------------------
b) If the price of a 30-pound watermelon is $4.20,
Use that fact to find "k":
4.2 = k(30)
k = 4.2/30
k = 2.1/15
k = 0.7/5
k = 7/50
So the Equation is P = (7/50)w
-------------------------------------------
then what is the price of a 20-pound watermelon?
P = (7/50)(20)
P = 14/5
Price = $2.80
==================
Cheers,
Stan H.